convert the raw score of 30 and 48 to a z-score


1
Expert's answer
2021-05-07T09:27:49-0400

Z score= (Xi-μ )/ δ 

Where μ is the mean

δ  is the standard deviation

Mean =30+48=78/2=39

Standard deviation= sqrt81=9

Z score of 30 = (30-39)/9

=-9/9= -1

Z score of 48 =(48-39)/9

= 9/9= 1